bill is three times as old as pedro. eighteen years from now, bill will be twice as old as pedro will be then. find the present age of both bill and pedro.
Answer by richwmiller(17219)

b=3p
b+18=2(p+18)
b+18=2p+36
b=2p+18
3p=2p+18
p=18
b=3*18=54
check
b+18=2(p+18)
54+18=2*(36)
72=72
ok